# Assessing the recombination scenario

### v1

@@ 4.1 $N=3$
For $N=3$, additional independent evidence against the recombination scenario is provided by the recent results of Refs. Henriksson et al. (2025a, b). These papers studied scaling dimensions of many operators in the Wilson-Fisher fixed point in $4-\epsilon$ expansion, at five loops. Suppose multiplet recombination occurs, so that $d=2+\epsilon$ and $d=4-\epsilon$ expansions are joined continuously. There is then an operator in the $4-\epsilon$ expansion in the same representation as the protected operator in $d=2+\epsilon$ expansion, whose scaling dimension reaches 2 for some $2<d<3$. The lightest such operator, for $N=3$, is De Cesare and Rychkov (2025)
@@
Results of Refs. Henriksson et al. (2025a, b) are exhaustive up to bare scaling dimension $6+O(\epsilon)$ and up to 2 Lorentz indices, and thus the five-loop anomalous dimension of (40) is available, see Eq.(15) of Henriksson et al. (2025a). In Fig. 2, we reproduce their result as a function of $2<d<4$, after Padé approximation. This suggests that the scaling dimension of $\widetilde{B}$ remains above 2, and hence this operator cannot connect continuously with the protected operator of the $2+\epsilon$ expansion. This contradiction disfavors recombination.
Figure 2: Scaling dimension of the operator $\widetilde{B}$ as a function of $d$ (dashed lines), obtained via a Padé approximation of the results of Henriksson et al. (2025a, b), with the use of Padém,n approximants. For this test, we show all approximants with $m+n=4$ and 5 ($m,n\geq 1$), discarding those which have poles in the range $2<d<4$. The Pade-approximated scaling dimension remains safely above that of the protected operator $B$ (solid line) for all the shown cases. We therefore conclude that the continuous connection between $d=2+\epsilon$ and $d=4-\epsilon$ expansions is unlikely.

### v2

@@ 4.1 $N=3$
It’s also interesting to probe the recombination scenario from the $d=4-\epsilon$ side. For $N=3$, the lightest operator in the $4-\epsilon$ expansion with the same Lorentz and global quantum numbers as the protected operator $B$ in $d=2+\epsilon$ expansion is De Cesare and Rychkov (2025)
@@
We can study the dimension of this operator as a function of $d$. We expect two possibilities. Either 1) the dimension of $\tilde{B}$ approaches 2 the dimension of $B$ as $d\to 2$, or 2) it crosses 2 at some $d=d_{c}>2$. The second possibility would allow recombination at $d=d_{c}$. The first possibility would correspond to the scenario in (De Cesare and Rychkov, 2025, Fig.5) that the O(3) WF and NLSM fixed points are two distinct families of CFTs which however share the main scaling dimensions in the limit $d\to 2$. [^10] The spectrum of O(3) NLSM at $d=2$ will be a subset of the limiting spectrum of the O(3) WF. The situation here will be analogous to that recently discussed by Zan for the Ising model Zan (2026), where the $d\to 2$ limiting spectrum is richer than the spectrum of the $d=2$ theory, to allow for the multiplet recombination as one moves away from $d=2$.
Recently, Refs. Henriksson et al. (2025a, b) studied scaling dimensions of many operators in the Wilson-Fisher fixed points in $4-\epsilon$ expansion, at five loops. Their results are exhaustive up to bare scaling dimension $6+O(\epsilon)$ and up to 2 Lorentz indices, and thus the five-loop anomalous dimension of (40) is available, see Eq.(15) of Henriksson et al. (2025a). [^11] We thank J. Henriksson for drawing our attention that there was a misprinting in this equation in the first arxiv version; here we are using the second arxiv version of Henriksson et al. (2025a).
In Fig. 2, we plot their result as a function of $2<d<4$, after Padé approximation. We see that the scaling dimension of $\widetilde{B}$ remains above 2 throughout most of the interval $2<d<4$, and approaches 2 at $d\approx 2.1$. This could in principle be interpreted as recombination at $d\approx 2.1$, but we find this unlikely: so close to $d=2$ our perturbative results precluding recombination should be trustworthy. The more natural interpretation of Fig. 2 is rather as an imperfect image of exact joining at $d=2$, imperfect since an approximate, five-loop result from the $d=4-\epsilon$ side is being used. In the future, this can be checked further by including more loops and seeing if the intersection point moves even closer to $d=2$.
Figure 2: Scaling dimension of the operator $\widetilde{B}$ as a function of $d$ (dashed lines), obtained via a Padé approximation of the results of Henriksson et al. (2025a, b), with the use of Padém,n approximants. For this test, we show all approximants with $m+n=4$ and 5 ($m,n\geq 1$), discarding those which have poles in the range $2<d<4$. The Pade-approximated scaling dimension remains safely above that of the protected operator $B$ for most of the interval, and reaches $\Delta=2$ for $d$ close to 2. We view this numerical evidence as supporting the fact that the NLSM and the WF CFT only coincide at $d=2$.